OpenStudy (anonymous):

Why is W not a subspace of the vector space? W is the set of all linear functions ax+b in c(continous functions)

OpenStudy (anonymous):

I dont know if that made sense but in simpler english W a linear function isnt a subspace of continous functions

OpenStudy (anonymous):

idk y cuz linear functions are continous so liek it is weird

OpenStudy (anonymous):

I think basically it wants you to figure out somehow how this linear function isnt a linear function either when multiplied by a scalar or when added to another function

OpenStudy (turingtest):

is it specifically the set of all functions of degree 1 ?

OpenStudy (turingtest):

...or is anything less that degree 1 allowed ?

OpenStudy (anonymous):

umm idk it says ax+b where a cant equal 0

OpenStudy (turingtest):

the only thing I can think is that this is specifically the set of all polynomials of degree exactly 1, and we may have the situation\[ax+b+(-ax)+c=b+c\]which would take it out if that space by lowering the degree if a and b can't be zero then we're done :D

OpenStudy (turingtest):

but that's different than what you said.... ax+b can't be 0 you say? what's that mean? lol

OpenStudy (anonymous):

no "a" can't equal 0

OpenStudy (turingtest):

oh, then we're done

OpenStudy (anonymous):

ok Thanks :D

hero (hero):

Look who's back :P

OpenStudy (anonymous):

Hey

hero (hero):

Howz it going?

OpenStudy (anonymous):

its going wish it wld be better but it is fine

hero (hero):

I'm already done with my course :P

OpenStudy (anonymous):

Oh well I am very far behind :D

hero (hero):

Yeah, next time, you should hang out with me when taking classes.

OpenStudy (anonymous):

Well i may be taking a break next term I may just take one course

OpenStudy (anonymous):

I dont really like linear algebra

OpenStudy (anonymous):

Calc is easy and fun it is more concrete while this is abstract for me

hero (hero):

lol

OpenStudy (anonymous):

Oh we r bothering turing Sorry

OpenStudy (turingtest):

Yeah I feel the same way only since I started using this site have I felt more comfortable with it the result of practice of course

OpenStudy (turingtest):

linear algebra I mean*

OpenStudy (anonymous):

I am gonna take abstract algebra one of these days and like i am nervous

OpenStudy (anonymous):

I know my brain wont be able to follow that

OpenStudy (turingtest):

I've never even looked at that myself...

OpenStudy (anonymous):

Turing I have another question

OpenStudy (phi):

Turing found this great page http://tutorial.math.lamar.edu/Classes/LinAlg/VectorSpaces.aspx The key fact for W not being a subspace is it says ax+b where a cant equal 0 to be a subspace, we need to satisfy a bunch of axioms, including There is a special object in V, denoted 0 and called the zero vector, such that for all u in V we have u+0=u so we need f(x)+ 0 = f(x) where 0 is in W. Of course, in this context, 0= 0x+0, which is not allowed because a cannot be 0. So W fails with this property

OpenStudy (anonymous):

So u r saying something diff than turing

OpenStudy (anonymous):

You r saying that a diff axiom fails

OpenStudy (phi):

You can fail for more than one reason. But when I went through the list given in the link, it looks like W fails with (e) and (f)

OpenStudy (anonymous):

okkkkk I see

OpenStudy (anonymous):

I have one more that i dont understand

OpenStudy (anonymous):

W is a set of all vectorsin R^3 whose compnents are pythagorean triple

OpenStudy (phi):

doesn't that fail for u+v is in W e.g. given u= (3,4,5), v= (5,12,13) is (8,16,18) in W?

OpenStudy (turingtest):

phi is saying the same thing more articulately

OpenStudy (anonymous):

kkk i get it

OpenStudy (anonymous):

ok Thanks guys :D U r awesome